OpenStudy (anonymous):

math 12 geometric sequences. which term has the value of 7over1024 in the geometric sequence 28,14,7

OpenStudy (mimi_x3):

It's like before; find the common ratio first.. \[\large \frac{T_{2}}{T_{1}} = \frac{14}{28} = \frac{1}{2} \] \[\large \frac{T_{3}}{T_{2}} = \frac{7}{14}=\frac{1}{2} \] Then use the formula.. \[\large T_n = ar ^{n-1}\] where a is the first term of the sequence and r is the common ratio.. \[\large T_n = \frac{7}{1024} \] \[\large ar ^{n-1} = \frac{7}{1024} \]

OpenStudy (anonymous):

yes I did that and i got 7 over 1024 = 28(1over2)to the power of n - 1

OpenStudy (anonymous):

and i divided each sides by 28

OpenStudy (anonymous):

and i used log method and I am not getting a right answer..

OpenStudy (dumbcow):

\[(n-1) \log(\frac{1}{2}) = \log(\frac{7}{1024*28})\] \[n-1 = 12\] n = 13

OpenStudy (anonymous):

ohhhhh okay thanks!!! :)